# Augmenting Subspace Optimization Methods with Linear Bandits

### v1

@@ 2 Preliminaries
To motivate our development, we begin with a lemma concerning a subspace gradient descent method defined via the iteration (3).
@@
Let 1 hold. Suppose $\alpha_{k}=\frac{1}{L}$, where $L$ is from 1. For all $k=0,1,\dots,K$, the sequence of $\{x_{k}\}$ generated by the iteration (3) satisfies
$f(x_{k})-f(x_{k+1})\geq\frac{1}{2L}\|(S_{k}^{\top}S_{k})^{-\frac{1}{2}}S_{k}^{\top}\nabla f(x_{k})\|^{2}.$
@@
By standard Taylor error arguments and using 1, we have
$\begin{array}[]{rl}f(x_{k+1})&\leq f(x_{k})+\langle\nabla f(x_{k}),x_{k+1}-x_{k}\rangle+\frac{L}{2}\|x_{k+1}-x_{k}\|^{2}\\ &=f(x_{k})+\langle\nabla f(x_{k}),-\frac{1}{L}P_{k}\nabla f(x_{k})\rangle+\frac{L}{2}\|-\frac{1}{L}P_{k}\nabla f(x_{k})\|^{2}\\ &=f(x_{k})-\frac{1}{L}\nabla f(x_{k})^{\top}S_{k}(S_{k}^{\top}S_{k})^{-1}S_{k}^{\top}\nabla f(x_{k})+\frac{1}{2L}\nabla f(x_{k})^{\top}S_{k}(S_{k}^{\top}S_{k})^{-1}S_{k}^{\top}S_{k}(S_{k}^{\top}S_{k})^{-1}S_{k}^{\top}\nabla f(x_{k})\\ &=f(x_{k})-\frac{1}{2L}\nabla f(x_{k})^{\top}S_{k}(S_{k}^{\top}S_{k})^{-1}S_{k}^{\top}\nabla f(x_{k})\\ &=f(x_{k})-\frac{1}{2L}\|(S_{k}^{\top}S_{k})^{-\frac{1}{2}}S_{k}^{\top}\nabla f(x_{k})\|^{2},\end{array}$
as intended. ∎
Notice that if $S_{k}=I_{d}$ for all $k$, then Lemma 1 recovers a well-known bound on function decrease for an iteration of gradient descent without any assumptions on the convexity of $f$. The next lemma is not about any particular optimization method but is an observation about a general $S_{k}$.
@@
For any $S_{k}\in\mathbb{R}^{d\times p_{k}}$ with full column rank,
$\|(S_{k}^{\top}S_{k})^{-\frac{1}{2}}S_{k}^{\top}\nabla f(x_{k})\|^{2}\leq\|\nabla f(x_{k})\|^{2},$
with equality attained if $\nabla f(x_{k})\in\text{Span}(S_{k})$. Here, we say $z\in\text{Span}(S_{k})$ provided there exists $y\in\mathbb{R}^{p_{k}}$ such that $S_{k}y=z.$
@@
Because $P_{k}=S_{k}(S_{k}^{\top}S_{k})^{-1}S_{k}^{\top}$ is a projection matrix, its eigenvalues are all either 0 or 1. Therefore, $I_{d}-S_{k}(S_{k}^{\top}S_{k})^{-1}S_{k}^{\top}\succeq 0$, and so
$\|\nabla f(x_{k})\|^{2}-\|(S_{k}^{\top}S_{k})^{-\frac{1}{2}}S_{k}^{\top}\nabla f(x_{k})\|^{2}=\nabla f(x_{k})^{\top}(I_{d}-S_{k}(S_{k}^{\top}S_{k})^{-1}S_{k}^{\top})\nabla f(x_{k})\geq 0,$
@@
Now, suppose $\nabla f(x_{k})\in\text{Span}(S_{k})$. Then, without loss of generality, we may assume that the columns of $S_{k}$ are $\nabla f(x_{k})/\|\nabla f(x_{k})\|$ and a set of $p_{k}-1$ unit vectors in $\mathbb{R}^{d}$, mutually orthogonal with $\nabla f(x_{k})$. Then,
$\|(S_{k}^{\top}S_{k})^{-\frac{1}{2}}S_{k}^{\top}\nabla f(x_{k})\|^{2}=\|S_{k}^{\top}\nabla f(x_{k})\|^{2}=\left|\displaystyle\frac{\nabla f(x_{k})}{\|\nabla f(x_{k})\|}^{\top}\nabla f(x_{k})\right|^{2}=\|\nabla f(x_{k})\|^{2}.$
Combining Lemma 1 and Lemma 2, we see that the lower bound on function decrease guaranteed by (3) proven in Lemma 1 is maximized when $S_{k}$ is well aligned with $\nabla f(x_{k})$ in the sense that (2) is satisfied, that is, $\|(S_{k}^{\top}S_{k})^{-\frac{1}{2}}S_{k}^{\top}\nabla f(x_{k})\|^{2}\approx\|\nabla f(x_{k})\|^{2}$. It is precisely this sort of alignment that appears in multiple analyses of methods in Framework 1. In fact, [9, 12, 18] all impose versions of the following two assumptions, which capture precisely this intuition of seeking to minimize the lower bound in Lemma 1.
@@
Let $\delta>0$ and $q>0$ be given. The sequence $\{S_{k}\}$ is generated in such a way that in the $k$th iteration, the inequality
$\|S_{k}^{\top}\nabla f(x_{k})\|^{2}\geq q\|\nabla f(x_{k})\|^{2}$ (4)
holds with probability at least $1-\delta$. When the event (4) is realized, we say that $S_{k}$ is $q$-well-aligned with $\nabla f(x_{k})$.
@@
There exists $S_{\max}>0$ such that the largest eigenvalue of $S_{k}^{\top}S_{k}$, $\lambda_{\max}(S_{k}^{\top}S_{k})$, satisfies $\lambda_{\max}(S_{k}^{\top}S_{k})>S_{\max}$ for all $k$.
@@
    $S_{k}$ is a Gaussian matrix; in other words, each entry of $S_{k}$ is independently drawn from $\mathcal{N}(0,1/p_{k})$. In this case, given both $\delta$ and $q$, 2 will hold provided $p_{k}\in\Omega\left(\frac{|\log(\delta)|}{(1-q)^{2}}\right)$. One can use matrix Bernstein inequality to derive high probability bounds to put $\lambda_{\max}(S_{k}^{\top}S_{k})\in\mathcal{O}(d/p_{k})$. Thus, while Gaussian sampling cannot guarantee that 3 holds, we can claim via an union bound over all $K$ matrices that 3 will hold with high probability. For more details on these fairly well-known results, see, for example, [34][Section 8.6].
@@
    $S_{k}$ is an $h_{k}$-hashing matrix; in other words, $S_{k}$ has exactly $h_{k}$ nonzero entries in each column, with the indices sampled independently, and each nonzero entry has value ±$\pm 1/\sqrt{h_{k}}$, each value being selected independently with probability $1/2$. In this case, given $\delta$ and $q$, then provided $h_{k}\in\Theta\left(\frac{|\log(\delta)|}{(1-q)}\right)$ and $p_{k}\in\Omega\left(\frac{|\log(\delta)|}{(1-q)^{2}}\right)$, 2 is satisfied [28]. Coarsely, we can bound $\|S_{k}\|\leq\|S_{k}\|_{F}=\sqrt{d}$, and so 3 is also satisfied with $S_{\max}=d$
@@
    $S_{k}$ is the first $p_{k}$ columns of a Haar-distributed orthogonal matrix, scaled by a factor $\sqrt{d/p_{k}}$. In this case, given $p_{k}$ and $q$, $S_{k}$ will satisfy 2 provided $\delta$ is the value of the cumulative distribution function of a beta distribution with shape parameters $\alpha=p_{k}/2$ and $\beta=(d-p_{k})/2$, evaluated at $q^{2}p_{k}/d$; see [30]. We also immediately obtain that 3 is satisfied with $S_{\max}=d/p_{k}$.
Remarkably, all three of these sketching techniques superficially allow for the subspace dimension $p_{k}$ to be independent of $d$. This dimension independence is less obvious with the Haar distribution, but numerical evidence suggests this is the case [30].
We will later impose 2 and 3 as minimal assumptions on a random sequence of sketches $\{S_{k}\}$ when considering the specific use of the subspace gradient descent method (3) as an optimization algorithm within Framework 1. With the goal in mind of satisfying something resembling (2) on every iteration of an arbitrary algorithm in Framework 1, we now state our bandit model in Framework 2.
1 Input: Initial point $x_{0}$, number of rounds $K$, function $f$ satisfying 1.
@@
3       Endogenous Action: Decision-maker chooses $s_{k}\in\mathbb{R}^{d}$.
4       Exogenous Action: Environment chooses $p_{k}\in[0,d]$ and $S_{k}\in\mathbb{R}^{d\times p_{k}}.$
5       Feedback: Decision-maker receives (an approximation of) $[S_{k},s_{k}]^{\top}\nabla f(x_{k})$.
6       State change: $x_{k}$ is updated according to an algorithm in Framework 1, e.g., (3).
@@
Framework 2 Description of linear bandit model
Framework 2 defines a game consisting of $K$ rounds. In the $k$th round of the game, the decision-maker must select one linear measurement $s_{k}\in\mathbb{R}^{d}$; we refer to this one-dimensional measurement $s_{k}$ as the decision-maker’s action from the action space $\mathbb{R}^{d}$. The decision-maker then receives directional derivatives in the directions defined by $S_{k}$ and the selected $s_{k}$; we assume the $p_{k}$ columns of $S_{k}$ are exogenously determined, for example, by a randomized sketching method. The environment then drifts slightly, in that the $k$th iteration of an optimization algorithm, for example, (3), is performed.
Because $x_{k}$, the state, changes as a result of performing each iteration of an optimization algorithm from Framework 1, the environment is nonstationary. Moreover, because we intend to solve an optimization problem (1) only once, it is inappropriate to view Framework 2 as a setting amenable to reinforcement learning; however, we do remark that if one is solving many similar optimization problems (1) via subspace optimization methods, learning a policy for playing Framework 2 could be interesting, but it is not the subject of this manuscript. As a result, it is most appropriate to analyze any method for playing the game in Framework 2 from the perspective of dynamic regret [7, 8]. We are particularly interested in dynamic regret in slowly varying or drifting environments [14, 15]. We say the environment is slowly varying because the reward function changes continuously with $x_{k}$ due to 1, and the state change is defined by an optimization iteration that seeks to asymptotically steer $\nabla f(x_{k})\to 0$.
@@
$R_{k}=\displaystyle\max_{S\in\mathbb{R}^{d\times p_{k}}}\|(S^{\top}S)^{-\frac{1}{2}}S^{\top}\nabla f(x_{k})\|-\|(S_{k}^{\top}S_{k})^{-\frac{1}{2}}S_{k}^{\top}\nabla f(x_{k})\|.$ (5)
As is standard in the definition of regret, the quantity $R_{k}$ in (5) is the gap between the best reward the decision-maker could have received and the reward that is realized. By Lemma 2, if the decision-maker chooses the action $s_{k}=\nabla f(x_{k})$, then the first term is maximized; and so we can simplify (5) to
$R_{k}=\|\nabla f(x_{k})\|-\|(S_{k}^{\top}S_{k})^{-\frac{1}{2}}S_{k}^{\top}\nabla f(x_{k})\|=\|\nabla f(x_{k})\|_{I-S_{k}(S_{k}^{\top}S_{k})^{-1}S_{k}^{\top}}.$
Thus, the decision-maker’s dynamic regret over the $K$ rounds of this game is defined as
$D_{K}:=\displaystyle\sum_{k=1}^{K}R_{k}=\displaystyle\sum_{k=1}^{K}\|\nabla f(x_{k})\|_{I-S_{k}(S_{k}^{\top}S_{k})^{-1}S_{k}^{\top}},$

### v4

@@ 2 Preliminaries
To motivate our development, we begin with a lemma guaranteeing the sufficient decrease attained in each iteration of Algorithm 2.
@@
Let 1 hold, in particular let $L$ be a global Lipschitz constant for $\nabla f$. Let a sequence of full column rank matrices $\{\mathbf{S}_{k}\}$ be given. For all $k=0,1,\dots,K$, the sequence of $\{\mathbf{x}_{k}\}$ generated by Algorithm 2 satisfies
$f(\mathbf{x}_{k})-f(\mathbf{x}_{k+1})\geq\frac{2\beta-1}{2L}\|(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-\frac{1}{2}}\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})\|^{2},$
where $\beta$ is the backtracking parameter from Algorithm 2.
@@
Suppose the while loop beginning in Algorithm 2 of Algorithm 2 runs through $N:=~\lceil\log_{\beta}(\min\{\beta,1/L\})\rceil$ many iterations so that $\alpha_{k}\in(\beta/L,1/L]$. By standard Taylor error arguments and using 1, we have
$\begin{array}[]{rl}f(\mathbf{x}_{k+1})&\leq f(\mathbf{x}_{k})+\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{k+1}-\mathbf{x}_{k}\rangle+\frac{L}{2}\|\mathbf{x}_{k+1}-\mathbf{x}_{k}\|^{2}\\ &=f(\mathbf{x}_{k})+\langle\nabla f(\mathbf{x}_{k}),-\alpha_{k}\mathbf{P}_{k}\nabla f(\mathbf{x}_{k})\rangle+\frac{L}{2}\|-\alpha_{k}P_{k}\nabla f(\mathbf{x}_{k})\|^{2}\\ &=f(\mathbf{x}_{k})-\alpha_{k}\nabla f(\mathbf{x}_{k})^{\top}\mathbf{S}_{k}(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-1}\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})\\ &+\frac{L\alpha_{k}^{2}}{2}\nabla f(\mathbf{x}_{k})^{\top}\mathbf{S}_{k}(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-1}\mathbf{S}_{k}^{\top}\mathbf{S}_{k}(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-1}\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})\\ &\leq f(\mathbf{x}_{k})-\frac{\beta}{L}\nabla f(\mathbf{x}_{k})^{\top}\mathbf{S}_{k}(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-1}\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})\\ &+\frac{1}{2L}\nabla f(\mathbf{x}_{k})^{\top}\mathbf{S}_{k}(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-1}\mathbf{S}_{k}^{\top}\mathbf{S}_{k}(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-1}\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})\\ \color[rgb]{0,0,0}\hfil&=f(\mathbf{x}_{k})-\frac{2\beta-1}{2L}\nabla f(\mathbf{x}_{k})^{\top}\mathbf{S}_{k}(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-1}\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})\\ &=f(\mathbf{x}_{k})-\frac{2\beta-1}{2L}\|(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-\frac{1}{2}}\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})\|^{2}\end{array}$
Thus, if the while loop of Algorithm 2 terminates before $N$ iterations, then $\alpha_{k}>\beta/L>(2\beta-1)/2L$, and so
$f(\mathbf{x}_{k+1})\leq f(\mathbf{x}_{k})-\alpha_{k}\|\mathbf{P}_{k}\nabla f(\mathbf{x}_{k})\|^{2}\leq f(\mathbf{x}_{k})-\frac{2\beta-1}{2L}\|\mathbf{P}_{k}\nabla f(\mathbf{x}_{k})\|^{2},$
proving the claim. ∎
The next lemma is not about any particular optimization method but is an observation about a general $\mathbf{S}_{k}$. In particular, Lemma 2 demonstrates the intuitive result that the norm of any projection of $\nabla f(\mathbf{x}_{k})$ onto $\mathbf{S}_{k}$ is necessarily upper bounded by the norm of $\nabla f(\mathbf{x}_{k})$, and equality can only be attained if $\nabla f(\mathbf{x}_{k})$ lies in the span of $\mathbf{S}_{k}$.
@@
For any $\mathbf{S}_{k}\in\mathbb{R}^{d\times p_{k}}$ with full column rank,
$\|(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-\frac{1}{2}}\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})\|^{2}\leq\|\nabla f(\mathbf{x}_{k})\|^{2},$
with equality attained if $\nabla f(\mathbf{x}_{k})\in\mathrm{Span}(\mathbf{S}_{k})$. Here, we say $\mathbf{z}\in\mathrm{Span}(\mathbf{S}_{k})$ provided there exists $\mathbf{y}\in\mathbb{R}^{p_{k}}$ such that $\mathbf{S}_{k}\mathbf{y}=\mathbf{z}.$
@@
Because $\mathbf{P}_{k}=\mathbf{S}_{k}(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-1}\mathbf{S}_{k}^{\top}$ is a projection matrix, its eigenvalues are all either 0 or 1. Therefore, $\mathbf{I}_{d}-\mathbf{S}_{k}(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-1}\mathbf{S}_{k}^{\top}\succeq 0$, and so
$\|\nabla f(\mathbf{x}_{k})\|^{2}-\|(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-\frac{1}{2}}\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})\|^{2}=\nabla f(\mathbf{x}_{k})^{\top}(\mathbf{I}_{d}-\mathbf{S}_{k}(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-1}\mathbf{S}_{k}^{\top})\nabla f(\mathbf{x}_{k})\geq 0,$
@@
Now, suppose $\nabla f(\mathbf{x}_{k})\in\mathrm{Span}(\mathbf{S}_{k})$. Then, without loss of generality, we may assume that the columns of $\mathbf{S}_{k}$ are $\nabla f(\mathbf{x}_{k})/\|\nabla f(\mathbf{x}_{k})\|$ and a set of $p_{k}-1$ unit vectors in $\mathbb{R}^{d}$, mutually orthogonal with $\nabla f(\mathbf{x}_{k})$. Then,
$\|(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-\frac{1}{2}}\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})\|^{2}=\|\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})\|^{2}=\left|\displaystyle\frac{\nabla f(\mathbf{x}_{k})}{\|\nabla f(\mathbf{x}_{k})\|}^{\top}\nabla f(\mathbf{x}_{k})\right|^{2}=\|\nabla f(\mathbf{x}_{k})\|^{2}.$
Combining Lemma 1 and Lemma 2, we see that the lower bound on per-iteration function decrease guaranteed by Algorithm 2 proven in Lemma 1 is maximized when $\mathbf{S}_{k}$ is well-aligned with $\nabla f(\mathbf{x}_{k})$ in the sense that (2) is satisfied, that is, $\|(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-\frac{1}{2}}\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})\|^{2}\approx\|\nabla f(\mathbf{x}_{k})\|^{2}$. It is precisely this sort of alignment that appears in multiple analyses of methods in Framework 1. In fact, [9, 12, 21] all impose versions of the following two assumptions, which capture precisely this intuition of seeking to minimize the lower bound in Lemma 1.
@@
Let $\delta\in[0,1)$ and $q\in(0,1)$ be given. The sequence $\{\mathbf{S}_{k}\}$ is generated in such a way that in the $k$th iteration, the inequality
$\|\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})\|^{2}\geq q\|\nabla f(\mathbf{x}_{k})\|^{2}$ (3)
holds with probability at least $1-\delta$. When the event (3) is realized, we say that $\mathbf{S}_{k}$ is $q$-well-aligned with $\nabla f(\mathbf{x}_{k})$.
@@
There exists $S_{\max}>0$ such that the largest eigenvalue of $\mathbf{S}_{k}^{\top}\mathbf{S}_{k}$$\lambda_{\max}(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})$, satisfies $\lambda_{\max}(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})>S_{\max}$ for all $k$.
@@
    $\mathbf{S}_{k}$ is a Gaussian matrix; in other words, each entry of $\mathbf{S}_{k}$ is independently drawn from $\mathcal{N}(0,1/p_{k})$. In this case, given both $\delta$ and $q$, 2 will hold provided $p_{k}\in\Omega\left(\frac{|\log(\delta)|}{(1-q)^{2}}\right)$  [^2] Throughout this manuscript, we will define $\Omega(\cdot)$, $\Theta(\cdot)$, and $\mathcal{O}(\cdot)$, as in [33]. One can use matrix Bernstein inequality to derive high probability bounds to put $\lambda_{\max}(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})\in\mathcal{O}(d/p_{k})$. Thus, while Gaussian sampling cannot guarantee that 3 holds, we can claim via an union bound over all $K$ matrices that 3 will hold with high probability. For more details on these fairly well-known results, see, for example, [39, Section 8.6].
@@
    $\mathbf{S}_{k}$ is an $h_{k}$-hashing matrix; in other words, each column of $\mathbf{S}_{k}$ is instantiated as a vector of zeros, and then exactly $h_{k}$ entries are chosen at random without replacement. Then, , each nonzero entry is assigned value $\pm 1/\sqrt{h_{k}}$, each value being selected independently with probability $1/2$. In this case, given $\delta$ and $q$, then provided $h_{k}\in\Theta\left(\frac{|\log(\delta)|}{(1-q)}\right)$ and $p_{k}\in\Omega\left(\frac{|\log(\delta)|}{(1-q)^{2}}\right)$, 2 is satisfied [32]. Coarsely, we can bound $\|\mathbf{S}_{k}\|\leq\|\mathbf{S}_{k}\|_{F}=\sqrt{d}$, and so 3 is also satisfied with $S_{\max}=d$.
@@
    $\mathbf{S}_{k}$ is the first $p_{k}$ columns of a Haar-distributed orthogonal matrix, scaled by a factor $\sqrt{d/p_{k}}$. In this case, given $p_{k}$ and $q$, $\mathbf{S}_{k}$ will satisfy 2 provided $\delta$ is chosen as $B_{p_{k}/2,(d-p_{k})/2}(q^{2}p_{k}/d)$, where $B_{\alpha,\beta}$ denotes the cumulative distribution function of the beta distribution with shape parameters $\alpha$ and $\beta$; see [35]. We also immediately obtain that 3 is satisfied with $S_{\max}=d/p_{k}$.
Remarkably, all three of these sketching techniques superficially allow for the subspace dimension $p_{k}$ to be independent of $d$. This dimension independence is less obvious with the Haar distribution (dimension independence effectively means $B_{p_{k}/2,(d-p_{k})/2}(q^{2}p_{k}/d)\in\Theta(1)$ as a function of $d$), but numerical evidence suggests this is the case [35]. Dimension independence may instead be directly provable from, e.g., [40, Lemma 6.1].
We will later impose 2 and 3 as minimal assumptions on a random sequence of sketches $\{\mathbf{S}_{k}\}$ when considering the specific use of Algorithm 2 as an optimization algorithm within Framework 1. With the goal in mind of satisfying something resembling (2) on every iteration of an arbitrary algorithm in Framework 1, we now state our bandit model in Framework 2.
1Input: Initial point $\mathbf{x}_{0}$, number of rounds $K$, function $f$ satisfying 1.
@@
3 Endogenous Action: Decision-maker chooses $\mathbf{s}_{k}\in\mathbb{R}^{d}$.
4 Exogenous Action: Environment chooses $p_{k}\in[0,d]$ and $\mathbf{S}_{k}\in\mathbb{R}^{d\times p_{k}}.$
5 Feedback: Decision-maker receives (an approximation of) $\mathbf{S}_{k},\mathbf{s}_{k}]^{\top}\nabla f(\mathbf{x}_{k})$.
6 State change: $\mathbf{x}_{k}$ is updated according to an iteration of an algorithm in Framework 1, e.g., an iteration of Algorithm 2
@@
Framework 3 Description of linear bandit model
Framework 2 defines a game consisting of $K$ rounds. In the $k$th round of the game, the decision-maker must select one linear measurement $\mathbf{s}_{k}\in\mathbb{R}^{d}$; we refer to this one-dimensional measurement $\mathbf{s}_{k}$ as the decision-maker’s action from the action space $\mathbb{R}^{d}$. The decision-maker then receives directional derivatives in the directions defined by $\mathbf{S}_{k}$ and the selected $\mathbf{s}_{k}$; we assume the $p_{k}$ columns of $\mathbf{S}_{k}$ are exogenously determined, for example, by a randomized sketching method. The environment then drifts slightly, in that the $k$th iteration of an optimization algorithm, for example, Algorithm 2, is performed.
Because $\mathbf{x}_{k}$, the state, changes as a result of performing each iteration of an optimization algorithm from Framework 1, the environment is nonstationary. Moreover, because we intend to solve an optimization problem (1) only once, it is inappropriate to view Framework 2 as a setting amenable to reinforcement learning; however, we do remark that if one is solving many similar optimization problems (1) via subspace optimization methods, learning a policy for playing Framework 2 could be interesting, but it is not the subject of this manuscript. As a result, it is most appropriate to analyze any method for playing the game in Framework 2 from the perspective of dynamic regret [6, 7]. We are particularly interested in dynamic regret in slowly varying or drifting environments [15, 16]. We say the environment is slowly varying because the reward function changes continuously with $\mathbf{x}_{k}$ due to 1, and the state change is defined by an optimization iteration that seeks to asymptotically steer $\nabla f(\mathbf{x}_{k})\to\mathbf{0}$.
@@
$R_{k}=\displaystyle\max_{\mathbf{S}\in\mathbb{R}^{d\times p_{k}}}\|(\mathbf{S}^{\top}\mathbf{S})^{-\frac{1}{2}}\mathbf{S}^{\top}\nabla f(\mathbf{x}_{k})\|-\|(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-\frac{1}{2}}\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})\|.$ (4)
As is standard in the definition of regret, the quantity $R_{k}$ in (4) is the gap between the best reward the decision-maker could have received and the reward that is realized. By Lemma 2, if the decision-maker chooses the action $\mathbf{s}_{k}=\nabla f(\mathbf{x}_{k})$, then the first term is maximized; and so we can simplify (4) to
$R_{k}=\|\nabla f(\mathbf{x}_{k})\|-\|(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-\frac{1}{2}}\mathbf{S}_{k}^{\top}\nabla f(\mathbf{x}_{k})\|=\|\nabla f(\mathbf{x}_{k})\|_{\mathbf{I}-\mathbf{S}_{k}(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-1}\mathbf{S}_{k}^{\top}}.$
Throughout this manuscript, we use the notation $\|\cdot\|_{\mathbf{M}}$ to denote the Mahalanobis norm with respect to an arbitrary positive definite matrix $\mathbf{M}$, that is, $\|\mathbf{x}\|_{\mathbf{M}}=(\mathbf{x}^{\top}\mathbf{M}\mathbf{x})^{1/2}$. Thus, the decision-maker’s dynamic regret over the $K$ rounds of this game is defined as
$D_{K}:=\displaystyle\sum_{k=1}^{K}R_{k}=\displaystyle\sum_{k=1}^{K}\|\nabla f(\mathbf{x}_{k})\|_{\mathbf{I}-\mathbf{S}_{k}(\mathbf{S}_{k}^{\top}\mathbf{S}_{k})^{-1}\mathbf{S}_{k}^{\top}},$